# Investigating the Fermi-Hubbard model by the Tensor-Backflow method

### v2

@@ Abstract
Recently, a variational wave-function based on the tensor representation of backflow corrections has achieved state-of-the-art energy precision in solving Fermi-Hubbard-type models tensorbackflow . However, the Fermi-Hubbard model is very challenging to solve, and the validity of a method relies on investigating the ground state’s physical property. For simplicity, we name the tensor representation of backflow corrections as the Tensor-Backflow in this work. We apply the Tensor-Backflow method to investigate the Fermi-Hubbard model on two-dimensional lattices as large as 256 sites, under various interaction strengths $U$, electron fillings $n$ and boundary conditions. Energy precision can be further improved by considering more backflow terms, such as considering backflow terms from next-nearest-neighbours or from all sites. Energy extrapolations on 64-site lattices give competitive results to the gradient optimized fPEPS with the bond dimension as large as $D$=20. For cases of $n$=0.875 and $U$=8 on the $16\times 16$ lattice under open boundary condition, by considering nearest-neighbour backflow terms, obtained energy is only $4.5\times 10^{-4}$ higher than the state-of-the-art method fPEPS with the bond dimension $D$=20. For periodic boundary condition, the variational wave-function is not enforced on any prior symmetry, meanwhile linear stripe order is successfully obtained. Under the same filling and $U$=10,12, energies obtained from initializations with the obtained wave-function for $U$=8 are lower than that from direct optimizations, meanwhile energies are competitive to or even better than state-of-the-art tensor network methods. For cases of $n$=0.8 and 0.9375, results consistent with the phase diagram from AFQMC are obtained by direct optimizations.
@@ I Introduction
The Fermi-Hubbard model is a simple quantum lattice model for describing correlated electrons. Solving the Fermi-Hubbard model is important for understanding the mechanism of superconductivity Hubbard_SC1 ; Hubbard_SC2 ; Hubbard_SC3 . However, investigating the Fermi-Hubbard model on a finite-sized lattice is very challenging. Because states with different orders near the ground state have energies so similar that the optimization is difficult. For finite systems, the lattice size should be sufficiently large to reduce the boundary effect. Therefore, to identify the property of the ground state, numerical methods are required to be unbiased and scalable. To achieve this goal, many numerical methods have been developped and benchmarked Hubbard_benchmark ; Manybody_benchmark ; Hubbard_DMET ; Hubbard_DMRG ; Hubbard_NN1 ; Hubbard_NN2 ; Hubbard_TNS ; tJ_fPEPS ; Hubbard_fPEPS .
@@
For example, the density-matrix-renormalization-group (DMRG) is suitable for one-dimensional or quasi one-dimensional systems Hubbard_benchmark ; Hubbard_DMRG ; Hubbard_DMRG2 ; Hubbard_DMRG3 ; Hubbard_DMRG4 , but the accuracy is not satisfactory for two-dimensional systems. The Quantum-Monte-Carlo (QMC) has very high accuracy, but computational complexity is too high for systems within the “sign problem” QMC_sign . The projected-entangled-pair-state (PEPS) can achieve high accuracy for two-dimensional lattices after variational optimizations tJ_fPEPS ; Hubbard_fPEPS ; PEPS_liu , however the computational complexity is high especially for periodic boundary condition (PBC). For the infinite-PEPS (iPEPS) iPEPS and the density-matrix-embedding-theory (DMET) Hubbard_DMET , the size of the supercell introduces a prior condition on the ground state.
@@
The variational-monte-carlo (VMC) has no “sign problem”, however its effectiveness depends on the variational wave-function. Recently, the neural-network based variational wave-functions have been intensively benchmarked RBM_science ; CNN_J1J2_1 ; CNN_J1J2_2 ; CNN_J1J2_3 ; J1J2_SOTA1 ; J1J2_SOTA2 ; J1J2_SOTA3 ; Hubbard_NN1 ; Hubbard_NN2 . By the deep structure, neural-networks can capture long-range correlations or entanglements in the quantum state, thus achieving state representation ability beyond the area-law entanglement NN_represent1 ; NN_represent2 . However, the complex structure of neural-networks introduces difficulties for optimizations. For example, although the backflow corrections by neural-networks have obtained high accuracy in continuous systems NN_continuous1 ; NN_continuous2 , energy precision for the latticed Fermi-Hubbard model is not satisfactory on large lattices Hubbard_NN1 ; Hubbard_NN2 . Recently, the Tensor-Backflow method has achieved state-of-the-art energy precision at reasonable optimization costs for Fermi-Hubbard-type models tensorbackflow . The wave-function is based on backflow corrections backflow1 ; backflow2 , meanwhile the representation ability is significantly enhanced by a way to increase the variational parameter number via embedding independent variables into independent degrees of freedoms into a tensor.
@@
In this work, we use the Tensor-Backflow along with one Lanczos optimization lanczos to investigate the Fermi-Hubbard model under various electron fills $n$, interaction strengths $U$ and boundary conditions on lattices as large as 256 sites. Without prior enforcements on symmetries, energies obtained are competitive to or even better than current state-of-the-art results, meanwhile energy precision can be further improved by considering more backflow terms. For $n$=0.875 and $U$=8 on the $16\times 16$ lattice, ground state with a linear stripe order is obtained by a direct optimization. Periods for spin-density-wave (SDW) and charge-density-wave (CDW) are 16 and 8, respectively. For the same filling and $U$=10,12, achieved energies from the initialization with the wave-function obtained for $U$=8 are lower than those from direct optimzations. For $n=$0.8 and 0.9375, we compare moderate strengths of $U$ and obtain results consistent with the phase diagram from AFQMC AFQMC_phase_diagram ; AFQMC_n0.9375 .
@@
This paper is organized as follows: Sec.II introduces the Tensor-Backflow method as well as the VMC optimization and a Lanczos optimization. Computational complexity of the method is introduced. Subsection.III.1 presents energy comparisons respect to current state-of-the-art methods such as fPEPS, tensor networks and AFQMC. Subsection.III.2 presents results for $n$=0.875, including energy comparisons, ground state patterns, charge and spin structure factors. Subsection.III.3 presents results for $n$=0.8 and 0.9375, with energy results as well as charge and spin structure factors. Sec.IV concludes this paper.
@@ II.1 Wave-function definitions
Based on literature tensorbackflow , we proposed an unrestricted form of backflow corrections: the summation for backflow terms are not only on positions but also on spins. The fictitious orbital is constructed as
$\phi^{B}_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})=\phi_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})+\sum_{j}c_{ij}[\mathbf{S}]\sum_{\sigma_{j}=\pm 1}\phi_{k\sigma_{k}}(\mathbf{r}_{j,\sigma_{j}}).$ (1)
Eq.(1) has limited number of variational parameters thus is not sufficient for state representation. For the lattice model, independent parameters in Eq.(1) are discrete values, thus we consider each independent parameter in the formula as a degree of freedom in the tensor $g$. The tensor $g$ has the dimension tensorbackflow :
@@
where $M$ is the particle number, $N$ is the site number, $d$ is the degrees of freedom each site and $Q$ is the number of backflow terms including the particle itself. For example, $Q=5,N$ when considering backflow terms from nearest-neighbours and all sites, respectively. We consider coefficients $c_{ij}$ in Eq.(1) depend on local configurations $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$ instead of $|\mathbf{S}\rangle$ for simplicity backflow1 ; backflow2 . Thus, first and second $d$ in Eq.(2) are for $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$, respectively. The number of variational parameters is the multiplication of all degrees of freedoms in tensor $g$.
@@
$M_{ik}^{B}=\sum_{q}g[i,k,\mathbf{s}(\mathbf{r}_{i}),q,\mathbf{s}(\mathbf{r}_{q})],$ (4)
@@ II.2 Optimization method
The wave-function is first optimized by the VMC, then further optimized by a Lanczos step. The energy and the $\alpha$-th parameter’s gradient are evaluated through the Markov-Chain-Monte-Carlo (MCMC) process :
@@
[…]
The variational parameters are updated according to the gradient descent method. Here we only adopt the first-order gradient descent due to the low optimization difficulty of the tensor representation. Because of the limited MC sample number, we take the sign of the gradient and apply a constant step size $\delta$: $\alpha^{\prime}=\alpha-\delta\mathrm{sgn}(G^{\alpha})$. Such parameter updating scheme has been successfully used in optimizing high dimensional tensors like the PEPS .
@@ II.3 Computational complexity
For the Tensor-Backflow wave-function, the complexity for generating a coefficient $w(\mathbf{S})$ is $\mathcal{O}(N^{3}Q)$ based on Eq.(3) and Eq.(4), where $\mathcal{O}(N^{3})$ is from the Slater determinant and $\mathcal{O}(Q)$ is from the summation in Eq.(4). When considering backflow terms from nearest-neighbour sites, there is $Q$=5, meanwhile the complexity is $\mathcal{O}(N^{3})$. The maximal representation ability is achieved by considering backflow terms from all sites: $Q=N$, meanwhile the complexity is $\mathcal{O}(N^{4})$. For the wave-function after a Lanczos optimization, based on Eq.(6), the complexity is $\mathcal{O}(N^{4}Q)$ for generating a coefficient $w(\mathbf{S})$.
The interval between two MC samples is $N$ during the optimization, thus the total complexity for one VMC step is $\mathcal{O}(N^{4}Q)$. In VMC optimizations, sample number and step size are important for achieving fully convergences. To identify these parameters, we have tested two optimization strategies: (1)under a particular sample number in the magnitude of $\mathcal{O}(10^{4})$, firstly setting the step size as $\mathcal{O}(10^{-3})$ and then reducing it to $\mathcal{O}(10^{-4})$; (2)under the sample number $\mathcal{O}(10^{4})$, using the step size $\mathcal{O}(10^{-3})$, then increasing the sample number to $\mathcal{O}(10^{5})$, lastly reducing the step size to $\mathcal{O}(10^{-4})$. We found that converged energies achieved by both strategies only have differences on the fourth decimal place. For both strategies, $\mathcal{O}(1\times 10^{4})$ optimization steps are applied for fully convergence. In the Lanczos optimization, the complexity is $\mathcal{O}(N^{5}Q)$ due to the term $\langle\Psi_{p=0}|\hat{H}^{2}|\Psi_{p=0}\rangle$ when calculating the variance $\sigma_{p}^{2}$.
@@ III Numerical results
In this section, we present numerical results for the two-dimensional Fermi-Hubbard model achieved by the Tensor-Backflow method. To demonstrate the method’s high accuracy, we firstly present energy comparisons respect to previous state-of-the-art results.
@@
where $t$ and $t^{\prime}$ are hopping strengths for NN and NNN hoppings, respectively. $U$ is the strength of on-site interactions, $\hat{c}_{i\sigma}^{\dagger}(\hat{c}_{i\sigma})$ creates (destroys) a particle of spin $\sigma$ on the $i$-th site, and the particle number operator $\hat{n}_{i\sigma}=\hat{c}_{i\sigma}^{\dagger}\hat{c}_{i\sigma}$. We set $t$=1 through our investigations. Energy results achieved by wave-functions before and after a Lanczos optimization are evaluated with the MC sample number in the magnitude of $\mathcal{O}(4.5\times 10^{4})$.

### v3

@@ Abstract
Recently, a variational wave-function based on the tensor representation of backflow corrections has achieved state-of-the-art energy precision in solving Fermi-Hubbard-type models tensorbackflow . However, the Fermi-Hubbard model is very challenging to solve, and the validity of a method relies on investigating the ground state’s physical property. For simplicity, we name the tensor representation of backflow corrections as the Tensor-Backflow in this work. We apply the Tensor-Backflow method to investigate the Fermi-Hubbard model on two-dimensional lattices as large as 256 sites, under various interaction strengths $U$, electron fillings $n$, next-nearest-neighbor hopping strengths $t^{\prime}$ and boundary conditions. Instead of considering backflow terms from all sites, competitive results are achieved by considering nearest-neighbor or next-nearest-neighbor backflow terms. Meanwhile the variational wave-function is not enforced on geometric symmetries. When $t^{\prime}$=0, by considering nearest-neighbor backflow terms, linear stripe order is sucessfully obtained for the case of nn=0.875 and UU=8 on the $16\times 16$ lattice under periodoc boundary condition. For a similar case under open boundary condition, obtained energy is only $4.5\times 10^{-4}$ higher than the state-of-the-art method fPEPS with the bond dimension DD=20. Comparing to state-of-the-art neural network results, energies are competitive and relative errors are below $5\times 10^{-3}$. For cases of nn=0.8 and 0.9375, results consistent with the phase diagram from AFQMC are obtained by direct optimizations. When $t^{\prime}$=-0.2, by considering next-nearest-neighbor backflow terms, obtained energies are competitive or even lower than state-of-the-art neural network results. For example, obtained energy for nn=0.875, UU=8 on the $12\times 12$ lattice under PBC is $8.1\times 10^{-4}$ lower comparing to that from the neural network state. Therefore, the Tensor-Backflow method has strong representation abilities for the Fermi-Hubbard model.
@@ I Introduction
The Fermi-Hubbard model is a simple quantum lattice model for describing correlated electrons. Solving the Fermi-Hubbard model is important for understanding the mechanism of superconductivity Hubbard_SC1 ; Hubbard_SC2 ; Hubbard_SC3 . However, investigating the Fermi-Hubbard model on a finite-sized lattice is very challenging. Because states with different orders near the ground state have energies so similar that the optimization is easily to get trapped on a local minima. For finite systems, the lattice size should be sufficiently large to reduce the boundary effect. Therefore, to identify the property of the ground state, numerical methods are required to be unbiased and scalable. To achieve this goal, many numerical methods have been developed and benchmarked Hubbard_benchmark ; Manybody_benchmark ; Hubbard_DMET ; Hubbard_DMRG ; Hubbard_NN1 ; Hubbard_NN2 ; Hubbard_NN3 ; Hubbard_NN4 ; Hubbard_TNS ; tJ_fPEPS ; Hubbard_fPEPS .
@@
For example, the density-matrix-renormalization-group (DMRG) is suitable for one-dimensional or quasi one-dimensional systems Hubbard_benchmark ; Hubbard_DMRG ; Hubbard_DMRG2 ; Hubbard_DMRG3 ; Hubbard_DMRG4 , but the accuracy is not satisfactory for two-dimensional systems. The Quantum-Monte-Carlo (QMC) has very high accuracy, but computational complexity is too high for systems within the “sign problem” QMC_sign_1 ; QMC_sign_2 ; QMC_sign_3 . The projected-entangled-pair-state (PEPS) can achieve high accuracy for two-dimensional lattices after variational optimizations tJ_fPEPS ; Hubbard_fPEPS ; PEPS_liu , however the computational complexity is high especially for periodic boundary condition (PBC). For the infinite-PEPS (iPEPS) iPEPS and the density-matrix-embedding-theory (DMET) Hubbard_DMET , the size of the supercell introduces a prior condition on the ground state.
@@
The variational-monte-carlo (VMC) has no “sign problem”, however its effectiveness depends on the variational wave-function. Recently, the neural-network based variational wave-functions have been intensively benchmarked RBM_science ; CNN_J1J2_1 ; CNN_J1J2_2 ; CNN_J1J2_3 ; J1J2_SOTA1 ; J1J2_SOTA2 ; J1J2_SOTA3 ; Hubbard_NN1 ; Hubbard_NN2 . By the deep structure, neural-networks can capture long-range correlations or entanglements in the quantum state, thus achieving state representation ability beyond the area-law entanglement NN_represent1 ; NN_represent2 . Although backflow corrections by neural-networks have obtained high accuracy in continuous systems NN_continuous1 ; NN_continuous2 ; NN_continuous3 For the lattice Fermi-Hubbard model, high optimization difficulties of neural networks lead to unsatisfied energy precision on large lattices Hubbard_NN1 ; Hubbard_NN2 or requirements of pre-conditions on wave-functions Hubbard_NN3 ; Hubbard_NN4 . Recently, the Tensor-Backflow method has achieved state-of-the-art energy precision at reasonable optimization costs for Fermi-Hubbard-type models tensorbackflow . The wave-function is based on backflow corrections backflow1 ; backflow2 , meanwhile the representation ability is significantly enhanced by the tensor representation of the wave-function.
@@
In this work, we use the Tensor-Backflow method well-developed in the previous literature tensorbackflow to investigate the Fermi-Hubbard model under various electron fillings $n$, interaction strengths $U$, next-nearest-neighbor hopping strengths $t^{\prime}$ and boundary conditions on lattices as large as 256 sites. Without prior enforcements on geometric symmetries, achieved energies are competitive or even better than current state-of-the-art results. When $t^{\prime}$=0, the ground state with a linear stripe order is achieved without applying pinning-fields for the case of nn=0.875 and UU=8 on the $16\times 16$ lattice under PBC. Periods for spin-density-wave (SDW) and charge-density-wave (CDW) are 16 and 8, respectively. For cases of nn=0.875, UU=10 and 12, linear stripe orders are preserved when transferring from the wave-function converged for UU=8, meanwhile energies are lower comparing to other orders achieved by direct optimzations. For cases of n=n=0.8 with UU=4, 8 and nn=0.9375 with UU=4, results are consistent with the phase diagram of AFQMC AFQMC_phase_diagram ; AFQMC_n0.9375 . For more challenging cases of $t^{\prime}$=-0.2 where non-zero superconductivity is expected Hubbard_SC1 , energy achieved by Tensor-Backflow is competitive or even lower than those from the state-of-the-art neural network state (NQS) Hubbard_NN3 , demonstrating the strong representation ability of Tensor-Backflow.
@@
This paper is organized as follows: Sec.II introduces the Tensor-Backflow method as well as the VMC optimization and a Lanczos optimization. Computational complexity of the method is introduced. Subsec.III.1 compares two optimization strategies. Subsec.III.2 presents energy comparisons respect to current state-of-the-art methods such as fPEPS, AFQMC, NQS, etc. Subsec.III.3 presents results for nn=0.875 with $t^{\prime}$=0 and $t^{\prime}$=-0.2, including energy comparisons, ground state patterns, charge and spin structure factors. Subsec.III.4 presents results for nn=0.8 and 0.9375, with energy results as well as charge and spin structure factors. Sec.IV concludes this paper.
@@ II.1 Wave-function definitions
In this subsection, we briefly introduce the Tensor-Backflow method developed in the previous literature tensorbackflow . An unrestricted form of backflow corrections is proposed in literature tensorbackflow : the summation for backflow terms are not only on positions but also on spins. The fictitious orbital is constructed as
$\phi^{B}_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})=\phi_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})+\sum_{j}c_{ij}[\mathbf{S}]\sum_{\sigma_{j}=\pm 1}\phi_{k\sigma_{k}}(\mathbf{r}_{j,\sigma_{j}}),$ (1)
where $c_{ij}[\mathbf{S}]$ is the coefficient depending on the configuration $|\mathbf{S}\rangle$ and $\phi_{k\sigma_{k}}$ is the Hatree-Fock (HF) orbital.
Based on the backflow formula, the NQS can improve the representation ability by learning backflow terms Hubbard_NN1 ; Hubbard_NN3 ; NN_continuous1 ; NN_continuous2 ; NN_continuous3 . Comparing to NQS Hubbard_NN2 ; Hubbard_NN3 , we use tensor representation of the wave-function to improve the representation ability, meanwhile keeping the optimization feasible for first-order-gradient methods tensorbackflow .
For a function with discrete variables of $x_{1}$ and $x_{2}$: $y=f(x_{1},x_{2})$, it can be represented by a tenesor $\xi$: $y=\xi_{x_{1},x_{2}}$, where $x_{1}$ and $x_{2}$ are indices in the $\xi$. We consider coefficients $c_{ij}$ in Eq.(1) depend on local configurations $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$ instead of $|\mathbf{S}\rangle$ for simplicity backflow1 ; backflow2 . Then, the independent variables in Eq.(1) are particle position $\textbf{r}_{i}$, orbital number $k$ for $\sigma_{k}=\pm 1$, backflow indices $j$, configurations $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$. Based on independent degrees of freedom in the wave-function, a tensor $g$ is constructed with each dimension in $g$ corresponding to the independent variable in Eq.(1) respectively. Therefore, the total dimension of $g$ is tensorbackflow :
@@
where $M$ is the site number, $N$ is the total particle number, $d$ is the degrees of freedom each site and $Q$ is the number of backflow terms including the particle itself. First and second $d$ are for $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$, respectively. When considering backflow terms from nearest-neighbors and all sites, there is: $Q=5$ and $N$, respectively.
@@
$M_{ik}^{B}=\sum_{q}g[\mathbf{r}_{i},k,\mathbf{s}(\mathbf{r}_{i}),q,\mathbf{s}(\mathbf{r}_{q})],$ (4)
@@
Therefore, the wave-function $w(\mathbf{S})$ is generated by a summation of elements from tensor $g$ based on the backflow rule. In the wave-function, there are no enforcements on translational or rotational symmetries, and the total number of variational parameters is the multiplication of all indices of $g$. Comparing to representing correlations by decomposed tensors CP_tensor , in this work, the representation ability is achieved by the tensor $g$ natually containing high order correlations tensorbackflow . For example, $g$ can represent high order correlations between $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$. Such flexible representation ability of $g$ as well as the simple construction of the wave-function lead to low optimization difficulty, therefore competitive precision is achieved within the first-order-gradient method.
@@ II.2 Optimization method
The wave-function is firstly optimized by the VMC, then further optimized by a Lanczos step. The energy and the $\alpha$-th parameter’s gradient are evaluated through the Markov-Chain-Monte-Carlo (MCMC) process Hubbard_fPEPS ; PEPS_liu :
@@
where the local energy is $E_{\mathrm{loc}}(\mathbf{S})=\sum_{\mathbf{S}^{\prime}}\frac{w(\mathbf{S}^{\prime})}{w(\mathbf{S})}\langle\mathbf{S}^{\prime}|\hat{H}|\mathbf{S}\rangle$, the $O_{\mathrm{loc}}(\mathbf{S})^{\alpha}=\frac{1}{w(\mathbf{S})}\frac{\partial w(\mathbf{S})}{\partial\alpha}$, and $\langle\cdots\rangle$ denotes the average on MCMC samples.
The variational parameters are updated according to the gradient descent method. Here we only adopt the first-order gradient descent due to the low optimization difficulty of the tensor representation. Because of the limited MC sample number, we take the sign of the gradient and apply a constant step size $\delta$: $\alpha^{\prime}=\alpha-\delta\mathrm{sgn}(G^{\alpha})$. Such parameter updating scheme has been successfully used in optimizing high dimensional tensors like the PEPS tJ_fPEPS ; PEPS_liu .
@@ II.3 Computational complexity
For the Tensor-Backflow wave-function, the complexity for generating a coefficient $w(\mathbf{S})$ is $\mathcal{O}(N^{3}Q)$ based on Eq.(3) and Eq.(4), where $\mathcal{O}(N^{3})$ is from the Slater determinant and $\mathcal{O}(Q)$ is from the summation in Eq.(4).
When considering backflow terms from nearest-neighbor (NN) sites, there is QQ=5. The representation ability can be improved by considering backflow terms in the range of next-nearest-neighbor (NNN) sites, there is QQ=9. We denote the above two types as NN and NNN backflows, and the complexities are $\mathcal{O}(N^{3})$ for both backflow types. The maximal representation ability is achieved by considering backflow terms from all sites: $Q=M$, meanwhile the complexity is $\mathcal{O}(MN^{3})$.
When generating a coefficient for the wave-function after a Lanczos optimization, the complexity is $\mathcal{O}(N^{4}Q)$ for generating a coefficient $w(\mathbf{S})$ based on Eq.(6). During the Lanczos optimization, because of the term $\langle\Psi_{p=0}|\hat{H}^{2}|\Psi_{p=0}\rangle$ when calculating the variance $\sigma_{p}^{2}$, the complexity is $\mathcal{O}(N^{5}Q)$.
@@ III Numerical results
In this section, we present numerical results for the two-dimensional Fermi-Hubbard model achieved by the Tensor-Backflow method. To demonstrate the method’s high accuracy, we firstly present energy comparisons respect to state-of-the-art results.
@@
where $t$ and $t^{\prime}$ are strengths for NN and NNN hoppings, respectively. $U$ is the strength of on-site interactions, $\hat{c}_{i\sigma}^{\dagger}(\hat{c}_{i\sigma})$ creates (destroys) a particle of spin $\sigma$ on the $i$-th site, and the particle number operator $\hat{n}_{i\sigma}=\hat{c}_{i\sigma}^{\dagger}\hat{c}_{i\sigma}$. We set tt=1 through our investigations. Energy results achieved by wave-functions before and after a Lanczos optimization are evaluated with the MC sample number in the magnitude of $\mathcal{O}(4.5\times 10^{4})$.
III.1 Optimization details
In VMC optimizations, sample number and step size are important for achieving fully convergences. To identify these parameters, we have tested two optimization strategies:
  • (1)
    Under the sample number in the magnitude of $\mathcal{O}(4.5\times 10^{4})$, firstly setting the step size as $2\times 10^{-3}$ and then reducing it to $2\times 10^{-4}$;
  • (2)
    Under the sample number in the magnitude of $\mathcal{O}(4.5\times 10^{4})$, firstly setting the step size as $2\times 10^{-3}$, then increasing the sample number to $\mathcal{O}(4.5\times 10^{5})$ and maintaining the step size, finally reducing the step size to $2\times 10^{-4}$.
Figure 1: Comparisons of two optimization strategies demonstrated by nn=0.875, UU=8 and $t^{\prime}$=0 on the $4\times 16$ lattice under OBC, with NN backflow in wave-functions. Both strategies start from the same initial state. For strategy 1, sample number is fixed as $\mathcal{O}(4.5\times 10^{4})$, the step size is reduced from $2\times 10^{-3}$ to $2\times 10^{-4}$ at the 6000-th step. For strategy 2, sample number is increased from $\mathcal{O}(4.5\times 10^{4})$ to $\mathcal{O}(4.5\times 10^{5})$ at around 3000-th step, meanwhile maintaining the step size to $2\times 10^{-3}$. After about another 2500 steps, the step size is reduced to $2\times 10^{-4}$. Strategy 1 and 2 converge to -0.6773 and -0.6774, respectively.
Comparisons of two optimization strategies are depicted in Fig.(1). The case in the figure is nn=0.875, UU=8, $t^{\prime}$=0 on the $4\times 16$ lattice under open boundary condition (OBC), with NN backflow terms in the wave-function. From the figure, both strategies converge to similar energies. However, strategy 2 requires less optimization steps. Both strategies achieve similar ground state patterns after fully convergences, and comparisons of ground state patterns are detailed in Appendix.(A). Furthermore, we find that strategy 2 is beneficial to the convergence of all-site backflow, due to the much larger MC sample number. Therefore, strategy 2 is used for optimizing all-site backflow wave-functions. On lattices larger than 64 sites, for both optimization strategies, no more than $2\times 10^{4}$ optimization steps are used for fully convergence.